# Learning Continually by Spectral Regularization

### v1

@@ Abstract
footnotetext: Department of Computing Science, University of Alberta
Department of Computer Science, Stanford University
Google DeepMind
Canada CIFAR AI Chair
Loss of plasticity is a phenomenon where neural networks become more difficult to train during the course of learning. Continual learning algorithms seek to mitigate this effect by sustaining good predictive performance while maintaining network trainability. We develop new techniques for improving continual learning by first reconsidering how initialization can ensure trainability during early phases of learning. From this perspective, we derive new regularization strategies for continual learning that ensure beneficial initialization properties are better maintained throughout training. In particular, we investigate two new regularization techniques for continual learning: (i) Wasserstein regularization toward the initial weight distribution, which is less restrictive than regularizing toward initial weights; and (ii) regularizing weight matrix singular values, which directly ensures gradient diversity is maintained throughout training. We present an experimental analysis that shows these alternative regularizers can improve continual learning performance across a range of supervised learning tasks and model architectures. The alternative regularizers prove to be less sensitive to hyperparameters while demonstrating better training in individual tasks, sustaining trainability as new tasks arrive, and achieving better generalization performance.
@@ 1 Introduction
A longstanding goal of machine learning research is to develop algorithms that can learn continually and cope with unforeseen changes in the data distribution [1, 2]. Current learning algorithms, however, struggle to learn from dynamically changing targets and are unable to adapt gracefully to unforeseen changes in the distribution during the learning process [3, 4, 5, 6]. Such limitations can be seen as a byproduct of assuming, one way or another, the problem is stationary. Recently, there has been growing recognition of the fact that there are limitations to what can be learned from a fixed and unchanging dataset [7], that there are implicit non-stationarities in many problems of interest [8], and that some real-world problems benefit from learning continually [9, 10].
@@
The concept of plasticity has been receiving growing attention in the continual learning literature, where the loss of plasticity—a reduction in the trainability of the neural network [11, 5, 12]—has been noted as a critical shortcoming in current learning algorithms. Learning algorithms, and more specifically neural networks, that are performant in the non-continual learning setting often struggle when applied to continual learning problems. Settings where a neural network must continue to learn after changes occur in the observations exhibit a striking loss of plasticity such that learning slows down or even halts after successive changes in the learning distribution [13, 6, 14].
@@
Several aspects of a learning algorithm have been found to contribute to, or mitigate, loss of plasticity. Examples include the type of optimizer [6], the step-size [15, 16], the number of optimiser iterations [5], and the use of specific regularizers [11, 14, 17]. Such factors hint that there might be simpler underlying optimisation principles that govern the loss of plasticity. For example, the success of several methods that regularize neural networks towards properties of the initialization suggests that some of those properties mitigate loss of plasticity [11, 17, 14].
@@
Various properties of neural network weights at initialization have been shown to be associated with trainability. Previous analyses have demonstrated that the trainability of deep neural networks can be improved by ensuring that the variance of hidden activations and gradients remains uniform across the different layers of the neural network [18], which has the effect of keeping the average singular value of the layerwise Jacobians close to one. A stronger condition is dynamical isometry, where all singular values are close to one, or exactly one in the case of an orthogonal initialization [19, 20, 21]. Over the course of learning, the singular values deviate from their initialization and tend to grow over time [22], which can reduce trainability and impede continual learning.
@@
Given the effectiveness of initialization in ensuring the trainability of a neural network, continual learning algorithms may benefit from sustaining the relevant properties of the weights. This is the motivation of several methods that use the initialization either through regularization [17] or weight reinitialization [11, 23, 6]. Such methods leverage initialization explicitly, either by regularizing parameters towards their initial values or by resetting parameters by resampling them from the initialization distribution. In this paper, we seek to more directly address loss of trainability by sustaining the key properties present at initialization, rather than replicate the initialization explicitly.
@@
We investigate the key properties of the initialization that ensure trainability, how these properties are lost over the course of learning, and the effects this has on continual learning. Based on this analysis, we then introduce two regularizers that try to maintain the identified properties: we propose to use the Wasserstein distance to regularize the parameters towards the initialization distribution, which is less restrictive than regularizing towards initial weights, and we propose spectral regularization to control the maximum singular value of the weight matrix to directly address the deviation of the singular values from initialization over the course of learning. Our experiments show that spectral regularization is more performant and less sensitive to hyperparameters than other regularizers across datasets, nonstationarities, and architectures. While well-tuned regularizers are often able to mitigate loss of plasticity to a varying degree, learning continually with spectral regularization is robust, achieving high initial and sustained performance. Finally, we show that spectral regularization is also capable of improving generalization with both Vision Transformer [24] and ResNet-18 [25] on continual versions of SVHN2 [26], CIFAR10, and CIFAR100 [27].
@@ 2 Problem Setting
We investigate the trainability of learning algorithms, and specifically neural networks, in task-agnostic continual supervised learning. We assume the learning algorithm uses a variant of stochastic gradient descent to update a neural network with $L$ layers, defined recursively as $f_{\theta}(x):=h_{L}(x)$, where $h_{0}(x)=x$ is the input vector, $h_{l+1}(x)=\phi(\mathbf{W}_{l}h_{l}(x)+\mathbf{b}_{l})$ with an element-wise activation function $\phi$ and parameters $\theta=\{\theta_{L},\dots,\theta_{1}\}$ where $\theta_{l}=\{\mathbf{W}_{l},\mathbf{b}_{l}\}$ is the weight and bias parameter for layer $l$ [^1] In this notation we suppressed the presence of biases, which can effectively be included in the parameters. We assume that $f_{\theta}$ is trained over a sequence of tasks: the $\tau$th task is specified by a distribution $p_{\tau}$ over the observation-target pairs (denoted by $(x,y)$). For simplicity we assume that the tasks (data distribution) change periodically after every $T$ iterations, but we consider the task-agnostic setting, where the learning algorithm does not have access to information about the tasks except through the data that it samples. For each iteration in the task, $t\in\left((\tau-1)T,\tau T\right]$, the learning algorithm must optimise the neural network’s parameters to minimize the objective
$J_{\tau}(\theta)=\mathbb{E}_{(x,y)\sim p_{\tau}}\big{[}\ell(f_{\theta}(x),y)\big{]}$
@@
for some loss function $\ell$. In this paper we consider gradient-based methods to learn the weight parameters; the basic version of gradient descent would update the parameters at iteration $t$ as $\theta^{(t+1)}=\theta^{(t)}-\alpha\nabla_{\theta}J_{\tau(t)}(\theta)\big{|}_{\theta=\theta^{(t)}}$, where $\tau(t)=\lfloor t/T\rfloor$ denotes the current task number. In practice this is done by considering the empirical error instead of $J_{\tau(t)}$ computed from samples $(x_{t},y_{t})\sim p_{\tau}$ for $t\in\left((\tau-1)T,\tau T\right]$ generated from the task distribution, and usually some more involved optimization algorithm.
@@
The evaluation of the continual learning algorithm is performed at the end of a task [5]. Specifically, the learning algorithm is evaluated at times $t=\tau T$, after being given $T$ iterations to learn on task $\tau$. It is commonly assumed that each task is sampled independently and identically, and that each task is equally difficult, meaning that a neural network which is randomly initialized on each task is able to reach a similar objective value for any particular task [6, 12, 5].
@@
Loss of plasticity in the continual learning literature can refer to either loss of trainability [11, 5] or to loss of generalization [15]. Because trainability is a requirement for learning and generalization, we focus on loss of trainability, using it interchageably with loss of plasticity. Specifically, we use loss of trainability to refer to the phenomenon that the objective value, $J_{\tau}(\theta^{(\tau T)})$, increases as a function of the task $\tau$. Equivalently, the performance measures, such as accuracy, decrease with new tasks. Performance can also decrease, then plateau, and in this case the network has lost trainability.
@@ 3 Spectral Properties and Continual Trainability
A growing body of work has identified that neural networks trained on continually changing distributions exhibit loss of trainability, slowing down (or even preventing) learning on new tasks [11, 13, 6, 5]. At a high level, the reason behind this phenomenon is simple: the solution of one task becomes the initialization for learning on the next task, and if this initialization is not sufficiently good to enable learning in the new task, we face the aforementioned problem. To avoid this issue, we need to keep the network parameters within a region that can serve as good initialization at the start of a task. This approach comes with two challenges: (i) determining a suitable region for parameter initialization; (ii) ensuring the initialization region does not prevent effective learning in the tasks encountered.
It is easy to satisfy (i) and (ii) in isolation: using a standard initialization algorithm, such as those of [28, 29], and then keeping the parameters fixed at this value satisfies (i) but clearly does not satisfy (ii). Conversely, not restricting the parameter space while using a standard learning algorithm addresses (ii) but not (i). We seek a balance of the two requirements. Note that (i) and (ii) are not binary properties, hence there can be a trade-off (e.g., allowing less favorable parameters for next task initialization allow better optimization in the current task). It is worth noting that such an approach has been successfully applied in the context of online convex optimization in changing environments, where several algorithms can be recognized as running a standard learning algorithm (such as mirror descent) on a carefully selected subset of the parameter space (see, e.g., [30, 31, 32]).
Our approach is the following: to address problem (i), we first identify key properties that initialization algorithms [28, 29] impose on the starting parameters of a neural network. Then, we propose two regularizers targeting these properties strikes a good balance between (i) and (ii).
@@ 3.2 An illustrative example
To make these points more concrete, consider a single layer neural network $f_{\theta}(x)=\theta_{2}\texttt{ReLU}(\theta_{1}x)$ mapping $x\in\mathbb{R}^{2}$ to $\mathbb{R}$. Suppose that the first task is to fit $x=(1,0)^{\top}$ and $y=0$ in mean-squared error. An optimal solution with $f_{\theta}(x)=y$ is $\theta_{1}=\begin{bmatrix}1&0\\ 0&a\end{bmatrix}$ and $\theta_{2}=(0,a)$ for an arbitrarily small value of $a\geq 0$. Now consider that the next task is to fit $x=(0,1)^{\top}$ and $y=1$. The gradients of the loss $\ell(f_{\theta}(x),y)=\frac{1}{2}(y-\ell(f_{\theta}(x))^{2}$ are, for the given new $(x,y)$ pair, $\nabla_{\theta_{2}}\ell(f_{\theta}(x),y)=(f_{\theta}(x)-y)\texttt{ReLU}(\theta_{1}x)=(a^{2}-1)(0,a)^{\top}$ and $\nabla_{\theta_{1}}\ell(f_{\theta}(x),y)=(f_{\theta}(x)-y)x\texttt{ReLU}^{\prime}(\theta_{1}x)^{\top}=(a^{2}-1)\begin{bmatrix}0&0\\ 0&a\end{bmatrix}$. Performing updates with these gradients keeps the $\Theta(a)$ parameters at the same order while maintaining a loss of approximately 1. The condition number of the weight matrix $\theta_{1}$ is $\kappa=1/a$, which requires $\Theta(1/a)=\Theta(\kappa)$ update steps to make significant progress during learning. (A more detailed analysis of another example is given in Appendix A.2.)
@@ 3.3 Trainability and Effective Gradient Diversity
We can generalize the above observation to a problem of reduced gradient diversity. By gradient diversity, we mean the spread of singular vectors in the matrix of per-example stochastic gradients, $\mathbf{G}=[\mathbf{g}_{1},\ldots,\mathbf{g}_{m}]$. For this analysis, we focus on the rank, which is one particular summary statistic for the set of singular values that counts the number of non-zero singular values, $\text{rank}(\mathbf{G})=|\{i:\sigma_{i}(\mathbf{G})>0\}|$. However, for practical purposes the rank is problematic because it is unstable to perturbations [34, Theorem 1]. In our experiments below we will consider the condition number, $\nicefrac{{\sigma_{1}(\mathbf{G})}}{{\sigma_{m}(\mathbf{G})}}$ and effective rank, $\text{erank}(\mathbf{G})=\sum\bar{\sigma}_{i}(\mathbf{G})\log\bar{\sigma}_{i}(\mathbf{G})$ where $\bar{\sigma}_{i}(\mathbf{G})=\nicefrac{{\sigma_{i}(\mathbf{G})}}{{\sum_{i}\sigma_{i}(\mathbf{G})}}$ [35]. These measures are particularly sensitive to a growing largest singular value, if it grows faster than the smallest singular values (our experiments show this to be the case in Section 5.3). In this case, we say there is a reduction in the effective gradient diversity.
@@
A reduction in the rank of the gradient matrix results in collinear gradients on different datapoints, limiting the diversity of gradient directions used in the updates. This can have an adverse effect on learning. In the extreme case, where the gradient matrix is rank one, every datapoint provides a gradient in the same direction, even if the datapoints correspond to different classes. Consider the gradient of the loss on a particular datapoint $(x_{i},y_{i})$ with respect to the weight matrix of a hidden layer, $\theta_{l}\in\mathbb{R}^{d\times d}$, which can be written recursively as $\mathbf{G}_{l,i}=\nabla_{\theta_{l}}\ell(f_{\theta}(x_{i}),y_{i})=\delta_{l,i}h_{l-1,i}^{\top}$, where $\delta_{l,i}=\theta_{l+1}^{\top}\delta_{l+1,i}\mathbf{D}_{l,i}$ is the error gradient from the previous layer with $\delta_{L,i}=\nicefrac{{\partial\ell(f_{\theta}(x_{i}),y_{i})}}{{\partial f_{\theta}}}$ and $\mathbf{D}_{l,i}=\text{Diag}(\phi^{\prime}([\theta_{l}h_{l,i}]_{1}),\dotso,\phi^{\prime}([\theta_{l}h_{l,i}]_{d}))$. We can rewrite the gradient in terms of its vectorization $\mathbf{g}_{l,i}=\text{vec}(\mathbf{G}_{l,i})=(\mathbf{I}_{d}\otimes\theta_{l+1}^{\top})\mathbf{v}_{l,i}$ where $\mathbf{I}_{d}$ is the $d\times d$ identity matrix and $\mathbf{v}_{1,i}=\text{vec}(\delta_{l+1,i}\mathbf{D}_{l,i}h_{l-1,i}^{\top})\in\mathbb{R}^{d^{2}}$ is the vectorization of data-dependent terms. The matrix of gradients for different datapoints is the concatenation of the per-example gradients, $\mathbf{G}_{l}=[\mathbf{g}_{l,1},\mathbf{g}_{l,2},\dotso,\mathbf{g}_{l,m}]=(\mathbf{I}_{d}\otimes\theta_{l+1}^{\top})\mathbf{V}_{l}$ where $\mathbf{V}_{l}=[\mathbf{v}_{l,1},\mathbf{v}_{l,2},\dotso,\mathbf{v}_{l,m}]$. The rank of $\mathbf{G}_{l}$ is upper bounded as
@@
Thus, if the (effective) rank of $\theta_{l+1}$ decreases, then the (effective) rank of the gradient matrix may decrease as well. The rank of the gradient matrix can even decrease due to rank reduction in parameters at other layers, or through the rank of the representation, which others have noted to occasionally correlate with loss of trainability [6, 5, 17].
3.4 Spectral Properties Away From Initialization During Continual Learning
@@
It has been observed empirically that the parameter norm grows at a $\sqrt{t}$ rate in large-scale self-supervised learning [36]. Similar observations have been made in continual learning, provided that the neural network does not stop learning due to loss of trainability [13, 6, 37]. A growing parameter norm is problematic from an optimisation perspective. Specifically, the parameter norm under consideration is the Frobenius norm, equal to the sum of squared singular values, $\|\mathbf{\theta}\|_{F}^{2}=\sum_{i}\sum_{j}[\theta]_{ij}^{2}=\sum_{i}\sigma_{i}(\theta)^{2}$. Thus, a growing parameter Frobenius norm is equivalent to an increasing sum of the squared singular values. In particular, since $\|\theta_{l}\|_{F}\leq\sqrt{\text{rank}(\theta_{l})}\sigma_{max}(\theta_{l})$ where $\sigma_{max}(\theta_{l})$ denotes the spectral norm (defined as the largest singular value) [38], if the parameter (Frobenius) norm of layer $l$ grows at a $\sqrt{t}$ rate, then the spectral norm of the parameter matrix $\theta_{l}$ also increases at the same rate. This leads to an increase in the spectral norm of the layerwise Jacobian, $\sigma_{d}(\mathbf{D}_{l})\sigma_{max}(\theta_{l})\leq\sigma_{max}(\mathbf{J}_{l})$, which may harm trainability. Moreover, as we argued in the previous section, an increase in the spectral norm can result in reduced effective gradient diversity.
4 Regularizing for Trainability in Continual Learning

### v2

@@ Abstract
footnotetext: Correspondence to: Alex Lewandowski <lewandowski@ualberta.ca>.
Loss of plasticity is a phenomenon where neural networks can become more difficult to train over the course of learning. Continual learning algorithms seek to mitigate this effect by sustaining good performance while maintaining network trainability. We develop a new technique for improving continual learning inspired by the observation that the singular values of the neural network parameters at initialization are an important factor for trainability during early phases of learning. From this perspective, we derive a new spectral regularizer for continual learning that better sustains these beneficial initialization properties throughout training. In particular, the regularizer keeps the maximum singular value of each layer close to one. Spectral regularization directly ensures that gradient diversity is maintained throughout training, which promotes continual trainability, while minimally interfering with performance in a single task. We present an experimental analysis that shows how the proposed spectral regularizer can sustain trainability and performance across a range of model architectures in continual supervised and reinforcement learning settings. Spectral regularization is less sensitive to hyperparameters while demonstrating better training in individual tasks, sustaining trainability as new tasks arrive, and achieving better generalization performance.
@@ 1 Introduction
A longstanding goal of machine learning research is to develop algorithms that can learn continually and cope with unforeseen changes in the data distribution (Ring, 1994, Thrun, 1998). Current learning algorithms, however, struggle to learn from dynamically changing targets and are unable to adapt gracefully to unforeseen changes in the distribution during the learning process (Zilly et al., 2021, Abbas et al., 2023, Lyle et al., 2023, Dohare et al., 2024). Such limitations can be seen as a byproduct of assuming, one way or another, that the problem is stationary. Recently, there has been growing recognition of the fact that there are limitations to what can be learned from a fixed and unchanging dataset (Hoffmann et al., 2022), that there are implicit non-stationarities in many problems of interest (Igl et al., 2021), and that some real-world problems benefit from learning continually (Han et al., 2022, Janjua et al., 2023).
@@
The concept of plasticity has been receiving growing attention in the continual learning literature, where the loss of plasticity—either a reduction in a neural network’s ability to train (Dohare et al., 2021, Lyle et al., 2022, Elsayed and Mahmood, 2024), or ability to generalize (Ash and Adams, 2020, Zilly et al., 2021)—has been noted as a critical shortcoming in current learning algorithms. Learning algorithms, and more specifically neural networks, that are performant in the non-continual learning setting, often struggle when applied to continual learning problems. Settings where a neural network must continue to learn after changes occur in the data distribution exhibit a striking loss of plasticity such that learning slows down (Lyle et al., 2023) or even halts after successive changes (Abbas et al., 2023, Nikishin et al., 2022, Dohare et al., 2024).
@@
Several aspects of a learning algorithm have been found to contribute to or mitigate loss of plasticity. Examples include the type of optimizer (Dohare et al., 2024, Lyle et al., 2023), the step-size (Ash and Adams, 2020, Berariu et al., 2021), the number of optimiser iterations (Lyle et al., 2023), and the use of specific regularizers (Dohare et al., 2021, Lyle et al., 2022, Kumar et al., 2023, Lewandowski et al., 2023). Such factors hint that there might be simpler underlying optimisation principles that govern the loss of plasticity. For example, the success of several methods that regularize neural networks towards properties of the initialization suggests that some of those properties mitigate loss of plasticity (Dohare et al., 2021, Kumar et al., 2023, Lyle et al., 2022).
@@
The properties of neural network weights at initialization are associated with trainability in the early phases of learning. Previous analyses have demonstrated that the trainability of deep neural networks can be improved by ensuring that the initialization variance of hidden activations and gradients remains uniform across the different layers of the neural network (Glorot and Bengio, 2010), which has the effect of keeping the average singular value of the layerwise Jacobians close to one. A stronger condition is dynamical isometry, where all singular values are close to one, or exactly one in the case of an orthogonal initialization (Saxe et al., 2014, Pennington et al., 2017, Xiao et al., 2018). Over the course of learning, the singular values deviate from their initialization and tend to grow over time (Martin and Mahoney, 2021), which can reduce trainability and impede continual learning.
@@
Given the effectiveness of initialization in ensuring the trainability of a neural network, continual learning algorithms may benefit from sustaining the relevant properties of the weights. This is the motivation of several methods that use the initialization either through regularization (Kumar et al., 2023) or weight reinitialization (Dohare et al., 2021, Sokar et al., 2023, Dohare et al., 2024). Such methods leverage the initialization explicitly, either by regularizing parameters towards their initial values or by resetting parameters by resampling them from the initialization distribution. In this paper, we seek to more directly address the loss of trainability by sustaining the key properties present at initialization, thereby striking a better balance between trainability and performance compared to approaches that replicate the initialization explicitly.
@@
We investigate the key properties of the initialization that ensure trainability, how these properties are lost over the course of learning, and the effects this has on continual learning. We identify that deviations from the initial singular value distribution can result in low gradient diversity, thereby impeding continual learning. Based on this analysis, we then introduce a spectral regularizer to control the deviation of the singular values by keeping the maximum singular value of each layer close to one to directly address the deviation of the singular values from initialization over the course of learning. Our experiments show that spectral regularization is more performant and less sensitive to hyperparameters than other regularizers across datasets, nonstationarities, and architectures. While well-tuned regularizers are often able to mitigate loss of plasticity to a varying degree, learning continually with spectral regularization is robust, achieving high initial and sustained performance. In particular, we show that spectral regularization is also capable of improving generalization with both Vision Transformer (Dosovitskiy et al., 2021) and ResNet-18 (He et al., 2016) on continual versions of tiny-ImageNet (Le and Yang, 2015), CIFAR10, CIFAR100 (Krizhevsky, 2009), and SVHN2 (Netzer et al., 2011). Note that these datasets and architectures encompass all continual supervised learning experimental settings considered in the loss of plasticity literature. We also show that spectral regularization can improve the performance of soft-actor critic (Haarnoja et al., 2018) in reinforcement learning settings where loss of plasticity occurs due to primacy bias (Nikishin et al., 2022).
@@ 2 Problem Setting
We investigate the trainability of neural network learning algorithms in the task-agnostic setting. We denote a neural network with $L$ layers, defined recursively as $f_{\theta}(x):=h_{L}(x)$, where $h_{0}(x)=x$ is the input vector, $h_{l+1}(x)=\phi(\mathbf{W}_{l}h_{l}(x)+\mathbf{b}_{l})$ with an element-wise activation function, $\phi$, and parameters $\theta=\{\theta_{L},\dots,\theta_{1}\}$ where $\theta_{l}=\{\mathbf{W}_{l},\mathbf{b}_{l}\}$ is the weight and bias parameter for layer $l$ [^1] In this notation we suppressed the presence of biases, which can effectively be included in the parameters. We assume that $f_{\theta}$ is trained over a sequence of tasks: the $\tau$th task is specified by a distribution $p_{\tau}$ over the observation-target pairs, denoted by $(x,y)$. For simplicity we assume that the task (data distribution) changes periodically after every $T$ iterations. In addition, we consider the task-agnostic setting, where the learning algorithm does not have access to information about the task except through the data that it samples. For each iteration in the task, $t\in\left((\tau-1)T,\tau T\right]$, the learning algorithm optimises the neural network’s parameters to minimize the objective
$J_{\tau}(\theta)=\mathbb{E}_{(x,y)\sim p_{\tau}}\big{[}\ell(f_{\theta}(x),y)\big{]},$
@@
for some loss function $\ell$. In this paper, we consider gradient-based methods to learn the weight parameters; the basic version of gradient descent would update the parameters at iteration $t$ as $\theta^{(t+1)}=\theta^{(t)}-\alpha\nabla_{\theta}J_{\tau(t)}(\theta)\big{|}_{\theta=\theta^{(t)}}$, where $\tau(t)=\lfloor t/T\rfloor$ denotes the current task number. In practice this is done by considering the empirical error instead of $J_{\tau(t)}$ using samples from the task distribution, $(x_{t},y_{t})\sim p_{\tau}$, and usually some more involved optimization algorithm.
@@
The evaluation of the continual learning algorithm is performed at the end of a task (Lyle et al., 2023). Specifically, the learning algorithm is evaluated at times $t=\tau T$, after being given $T$ iterations to learn on task $\tau$. A useful assumption in some problems is that each task is sampled independently and identically, and that each task is equally difficult, meaning that a neural network which is randomly initialized on each task is able to reach a similar objective value for any particular task (Lyle et al., 2023, Elsayed and Mahmood, 2024, Dohare et al., 2024). While this assumption is useful, it is not always applicable. In class-incremental learning (Van de Ven et al., 2022), the addition of more classes increases the difficulty of subsequent tasks. Another example is the transient tasks induced over the course of reinforcement learning (Igl et al., 2021), which can have varying task difficulty and lead to loss of plasticity due to primacy bias (Nikishin et al., 2022). We also consider such settings here.
@@
Loss of plasticity in the continual learning literature can refer to either loss of trainability (Dohare et al., 2021, Lyle et al., 2023) or to loss of generalization (Ash and Adams, 2020). Because trainability is a requirement for learning and generalization, we focus primarily on loss of trainability. Specifically, we use loss of trainability to refer to the phenomenon that the objective value, $J_{\tau}(\theta^{(\tau T)})$, increases as a function of the task $\tau$. Equivalently, the performance measures, such as accuracy, decrease with new tasks. Under the assumption that the tasks are sampled independently and identically, this would suggest that the neural network’s trainability diminishes on new tasks.
@@ 3 Spectral Properties and Continual Trainability
At a high level, the reason behind loss of plasticity is simple: the solution of one task becomes the initialization for learning on the next task, and if this initialization is not sufficiently good to enable learning in the new task, we face the aforementioned problem. To avoid this issue, we need to keep the network parameters within a region that can serve as good initialization at the start of a task. This approach comes with two challenges: (i) determining a suitable region for parameter initialization on any given task; (ii) ensuring that training within the initialization region does not prevent effective learning on the current task.
It is easy to satisfy (i) and (ii) in isolation: using a standard initialization algorithm, such as those by Hinton and Salakhutdinov (2006), He et al. (2015), and then keeping the parameters fixed at this value satisfies (i) but clearly does not satisfy (ii). Conversely, not restricting the parameter space while using a standard learning algorithm addresses (ii) but not (i). We seek a balance of the two requirements. Note that (i) and (ii) are not binary properties; hence there can be a trade-off (e.g., the parameter that allows for better optimization on the current task may be less favourable as a parameter initialization for the next task). It is worth noting that such an approach has been successfully applied in the context of online convex optimization in changing environments, where several algorithms can be recognized as running a standard learning algorithm (such as mirror descent) on a carefully selected subset of the parameter space (e.g., Zinkevich, 2003, Herbster and Warmuth, 1998, György and Szepesvári, 2016).
In this paper, to address problem (i), we first identify key properties that initialization algorithms (Hinton and Salakhutdinov, 2006, He et al., 2015) impose on the starting parameters of a neural network. Then, we propose spectral regularization to sustain these properties while striking a good balance between (i) and (ii).
@@ 3.2 An illustrative example
To make these points more concrete, consider a single-layer neural network $f_{\theta}(x)=\theta_{2}\texttt{ReLU}(\theta_{1}x)$ mapping $x\in\mathbb{R}^{2}$ to $\mathbb{R}$. Suppose that the first task is to fit $x=(1,0)^{\top}$ and $y=0$ in mean-squared error. An optimal solution with $f_{\theta}(x)=y$ is $\theta_{1}=\begin{bmatrix}1&0\\ 0&a\end{bmatrix}$ and $\theta_{2}=(0,a)$ for an arbitrarily small value of $a\geq 0$. Now consider that the next task is to fit $x=(0,1)^{\top}$ and $y=1$. The gradients of the loss $\ell(f_{\theta}(x),y)=\frac{1}{2}(y-\ell(f_{\theta}(x))^{2}$ are, for the given new $(x,y)$ pair, $\nabla_{\theta_{2}}\ell(f_{\theta}(x),y)=(f_{\theta}(x)-y)\texttt{ReLU}(\theta_{1}x)=(a^{2}-1)(0,a)^{\top}$ and $\nabla_{\theta_{1}}\ell(f_{\theta}(x),y)=(f_{\theta}(x)-y)x\texttt{ReLU}^{\prime}(\theta_{1}x)^{\top}=(a^{2}-1)\begin{bmatrix}0&0\\ 0&a\end{bmatrix}$. Performing updates with these gradients keeps the $\Theta(a)$ parameters at the same order while maintaining a loss of approximately 1. The condition number of the weight matrix $\theta_{1}$ is $\kappa=1/a$, which requires $\Theta(1/a)=\Theta(\kappa)$ update steps to make significant progress during learning. (Another example showing this empirically is given in Appendix A.1.)
@@ 3.3 Trainability and Effective Gradient Diversity
We can generalize the above observation to a problem of reduced gradient diversity. By gradient diversity, we mean the spread of singular vectors in the matrix of per-example stochastic gradients, $\mathbf{G}=[\mathbf{g}_{1},\ldots,\mathbf{g}_{m}]$. If this matrix contains a few large singular values and many small singular values, then the gradients will be largely in the span of the singular vectors corresponding to the large singular values. For this analysis, we focus on the rank, which is one particular summary statistic for the set of singular values that counts the number of non-zero singular values, $\text{rank}(\mathbf{G})=|\{i:\sigma_{i}(\mathbf{G})>0\}|$. However, for practical purposes, the rank is problematic because it is unstable to perturbations (Feng et al., 2022, Theorem 1). In our experiments below we will consider the condition number, $\nicefrac{{\sigma_{1}(\mathbf{G})}}{{\sigma_{m}(\mathbf{G})}}$ and effective rank, $\text{erank}(\mathbf{G})=\sum\bar{\sigma}_{i}(\mathbf{G})\log\bar{\sigma}_{i}(\mathbf{G})$, where $\bar{\sigma}_{i}(\mathbf{G})=\nicefrac{{\sigma_{i}(\mathbf{G})}}{{\sum_{i}\sigma_{i}(\mathbf{G})}}$ (Roy and Vetterli, 2007). These measures make explicit the issues that arise if the largest singular value grows faster than the smallest singular values (our experiments show this to be the case in Section 3). In this case, the condition number increases and the effective rank decreases. We refer to this as a reduction in the effective gradient diversity.
@@
A reduction in the rank, or erank, of the gradient matrix results in collinear gradients on different datapoints, limiting the diversity of gradient directions used in the parameter update. This can have an adverse effect on learning. In the extreme case, where the gradient matrix is rank one, every datapoint provides a gradient in the same direction, even if the datapoints correspond to different classes. Consider the gradient of the loss on a particular datapoint $(x_{i},y_{i})$ with respect to the weight matrix of a hidden layer, $\theta_{l}\in\mathbb{R}^{d\times d}$, which can be written recursively for the parameters of layer $l$ as $\mathbf{G}_{l,i}=\nabla_{\theta_{l}}\ell(f_{\theta}(x_{i}),y_{i})=\delta_{l,i}h_{l-1,i}^{\top}$, where $\delta_{l,i}=\theta_{l+1}^{\top}\delta_{l+1,i}\mathbf{D}_{l,i}$ is the error gradient from the next layer with $\delta_{L,i}=\nicefrac{{\partial\ell(f_{\theta}(x_{i}),y_{i})}}{{\partial f_{\theta}}}$ and $\mathbf{D}_{l,i}=\text{Diag}(\phi^{\prime}([\theta_{l}h_{l,i}]_{1}),\dotso,\phi^{\prime}([\theta_{l}h_{l,i}]_{d}))$. We can rewrite the gradient in terms of its vectorization $\mathbf{g}_{l,i}=\text{vec}(\mathbf{G}_{l,i})=(\mathbf{I}_{d}\otimes\theta_{l+1}^{\top})\mathbf{v}_{l,i}$ where $\mathbf{I}_{d}$ is the $d\times d$ identity matrix and ${\mathbf{v}_{1,i}=\text{vec}(\delta_{l+1,i}\mathbf{D}_{l,i}h_{l-1,i}^{\top})\in\mathbb{R}^{d^{2}}}$ is the vectorization of data-dependent terms. The matrix of gradients for different datapoints is the concatenation of the per-example gradients, $\mathbf{G}_{l}=[\mathbf{g}_{l,1},\mathbf{g}_{l,2},\dotso,\mathbf{g}_{l,m}]=(\mathbf{I}_{d}\otimes\theta_{l+1}^{\top})\mathbf{V}_{l}$ where $\mathbf{V}_{l}=[\mathbf{v}_{l,1},\mathbf{v}_{l,2},\dotso,\mathbf{v}_{l,m}]$. The rank of $\mathbf{G}_{l}$ is upper bounded as
@@
Thus, if the (effective) rank of $\theta_{l+1}$ decreases, then the (effective) rank of the gradient matrix may decrease as well. The rank of the gradient matrix can even decrease due to rank reduction in parameters at other layers, or through the rank of the representation, which others have noted to occasionally correlate with loss of trainability (Lyle et al., 2023, Kumar et al., 2023, Dohare et al., 2024).
Why Do Spectral Properties Deviate From Initialization?
@@
In large-scale self-supervised learning, it has been observed empirically that the parameter norm grows at a rate of $\sqrt{t}$, where $t$ is the number of iterations (Merrill et al., 2021). Similar observations have been made in continual learning, provided that the neural network does not stop learning due to loss of trainability (Nikishin et al., 2022, Lyle et al., 2024, Dohare et al., 2024). A growing parameter norm is problematic from an optimization perspective. Specifically, the parameter norm that these works consider is the Frobenius norm, which is equal to the sum of squared singular values, $\|\mathbf{\theta}\|_{F}^{2}=\sum_{i}\sum_{j}[\theta]_{ij}^{2}=\sum_{i}\sigma_{i}(\theta)^{2}$. Thus, a growing parameter norm is equivalent to an increasing sum of the squared singular values. In particular, note that $\|\theta_{l}\|_{F}\leq\sqrt{\text{rank}(\theta_{l})}\sigma_{1}(\theta_{l})$ where $\sigma_{1}(\theta_{l})$ denotes the largest singular value, or the spectral norm (Golub and Van Loan, 2013). If the parameter norm of layer $l$ grows at a rate of $\sqrt{t}$, then the spectral norm of the parameter matrix $\theta_{l}$ also increases at the same rate. This leads to an increase in the spectral norm of the layerwise Jacobian, $\sigma_{d}(\mathbf{D}_{l})\sigma_{1}(\theta_{l})\leq\sigma_{1}(\mathbf{J}_{l})$, which can reduce effective gradient diversity and may harm trainability.
4 Spectral Regularization for Continual Learning